2)In a group of 20 adults, 4 out of 7 women and 2out of 13 men wear glasses. What is the probability that a person chosen at random from the group is a women or someone who wear glasses.
3) A firm of charted accountants has two vacancies for trainee students. In the past , 40% of the students who were offered the training contract have not reported to join.If two students are offered training contract, what is the probability that one will join?

Answer by stanbon(75887)

In a group of 20 adults, 4 out of 7 women and 2 out of 13 men wear glasses. What is the probability that a person chosen at random from the group is a women or someone who wear glasses.
---
P(w or g) = P(w) + P(g) - P(w and g)
= 7/20 + 6/20 - 4/20
= 9/20
=============================
3) A firm of charted accountants has two vacancies for trainee students. In the past , 40% of the students who were offered the training contract have not reported to join.If two students are offered training contract, what is the probability that one will join?
P(one will join) = 1 - (P(0 join) + P(both join))
= 1 - (0.6)^2 + (0.4)^2) = 1 - 0.36 - 0.14 = 1 - 0.5 = 0.5
====================================
Cheers,
Stan H.